# Benchmarking Foundation Models and Parameter-Efficient Fine-Tuning for Prognosis Prediction in Medical Imaging

### v1

@@ 4 Results and Discussion
In this section, we highlight the representative observations and outcomes obtained in the benchmark, obtained from fine-tuning the models on our set of datasets on COVID-19 prognosis tasks (both presented in presented in Section 3.2). All the results shown in the following sections are reported using the MCC score, which serves as evaluation metric for the results’ analysis. A complete summary of all results is provided in Table4 and Table 5 for the MCC and PR-AUC, where also standard deviations and the first and second best performing technique-model pairs are shown for each dataset. In this section, we present the findings emerging from benchmarking the models introduced in Section 3.2 on our suite of datasets. In the following discussion, we systematically examine different aspect and criticality raised in the introduction, structuring the analysis in two parts: a Setup paragraph, outlining the specific experimental setup (i.e., models and PEFT methods analyzed), followed by a Results paragraph, which presents and interprets the results obtained combining models, datasets, and fine-tuning strategies.
Figure 2: This figure includes only the fine-tuning techniques applicable to both $CNN$ and $FM$ architecture families. The mean performance over the test set folds for all fine-tuning methods is represented by the Refer to caption symbol, following the color scheme: $\mathbf{LoRA_{r=4}}$, $\mathbf{LoRA_{r=8}}$, $\mathbf{LoRA_{r=16}}$; $\mathbf{BitFit}$, $\mathbf{LP}$. The FFT method is uniquely represented by the Refer to caption symbol.
Figure 3: This figure displays the mean over all datasets performances respect only the fine-tuning techniques applicable to both $CNN$ and $FM$ architecture families. The overall mean for the fine-tuning methods is represented by the Refer to caption symbol, following the color scheme: $\mathbf{LoRA_{r=4}}$, $\mathbf{LoRA_{r=8}}$, $\mathbf{LoRA_{r=16}}$; $\mathbf{BitFit}$, $\mathbf{LP}$. The FFT method is uniquely represented by the Refer to caption symbol.
4.1.1 Setup.
We fine-tuned all the models introduced in Section 3.2 on the CXR prognosis tasks described in Section 3.1: to ensure a fair comparison from Table 3 we included only the fine-tuning techniques applicable to both architecture families which are AFCmr∈4,8,16, BitFit, LP, and LP.
4.1.2 Results.
Figure 2 presents the aggregated performance, obtained averaging the results across of the fine tuning techniques by dataset and by model.
We start our analysis with the smallest datasets, CAR and CC (Figures 2.c, 2.d), most PEFT methods fail to yield substantial performance gains under such constrained data conditions, with performance stabilizing around $20\%$ for CAR and $5\%$ for CC. However, certain FT approach combined with models show exceptions to this trend. Notably, BitFit and LoRA yield better results on DINOv2 variants, while FFT consistently outperforms other approaches on CNNs architectures, particularly DenseNet121 and ResNet18 for CC and three CNNs for CAR. These findings indicate that CNNs remain more resilient for extremely small datasets, likely due to their lower parameter count and the efficiency conferred by their inductive biases, even if not derived from pretraining the model on medical domain datasets. In contrast, FMs exhibit reduced performance under such conditions, primarily due to their larger parameterization, a limitation further exacerbated by the imbalance that characterize prognosis datasets. Despite this, the promising results achieved with PEFT on the smaller and base DINOv2 variants suggest that, with appropriate fine-tuning strategies, these models can still extract meaningful representations even in extreme low-data regimes. Nonetheless, it is evident that the DINOv2 large fails to generalize in this context—likely due to its higher parameter count, which is disproportionate to the limited dimensionality and scale of the available training data.
A second key insight emerges from comparing AFC and AFCm, two tasks that share the same data samples but differ in balancing conditions. In AFC (Figure 2.a), where the dataset is balanced, PEFT fine-tuning performs consistently across models regardless of the underlying architecture. At the same time DINOv2s and PubMedCLIP achieve the best performance, along with CNNs fine-tuned using FFT. This suggests that in scenarios with balanced data, PEFT techniques can successfully leverage pretrained representations without being strongly influenced by model choice [52]. However, in AFCm (Figure 2.b), where class distribution is highly imbalanced, PEFT performance drops significantly, yielding results around $10\%$ in MCC. Interestingly, PEFT methods remain effective only on DINOv2 variants, while FFT consistently outperforms other approaches across all models. These results confirm the strong sensitivity of PEFT techniques to data imbalance, where models struggle to generalize effectively under skewed class distributions.
Lastly, a comparative analysis between NYsmall and NYall (Figures 2.e, 2.f) highlights dataset size as a critical factor in PEFT effectiveness. In NYsmall, most PEFT methods fail to achieve meaningful improvements, with only certain DINOv2s variants showing moderate success, while FFT on PubMedCLIP emerges as the top performer. In NYall, PEFT techniques exhibit significantly higher performance, in some cases matching or even competing with FFT, particularly on DINOv2 and CLIP-Large. This confirms that PEFT methods require a sufficiently large dataset to effectively adapt and optimize their representations.
In conclusion, our findings further suggest that PEFT methods are highly sensitive to class imbalance and dataset size. Severe imbalance and small sample sizes negatively impact their performance, whereas larger datasets, even with moderate imbalance, result in significant improvements. DINOv2s exhibit an interesting trend in FFT setting: smaller variants tend to outperform larger ones, as seen in Figure 3. This suggests that excessive model complexity may not always translate to better fine-tuning outcomes in low-resource settings. Furthermore, in the comparison between MedCLIPc and its supervised pretrained equivalent ResNet50, it is evident that despite MedCLIPc was pretrained using SSL, the performance gains are lower than expected. Across most fine-tuning techniques, it achieves only a marginal improvement, with increases of just a few percentage points in the metric score.
Given our demonstration of the consistent benefits provided by PEFT applied to foundation models FMs in medical imaging, we henceforth concentrate our analysis on this paradigm. This decision aligns with established knowledge in the literature, notably Dutt et al. [51], who demonstrated that the advantages of PEFT typically grow with decreasing data volume and increasing model size, particularly benefiting small- to medium-sized datasets common in medical imaging. However, our analysis extends existing understanding by highlighting important limitations: we show that CNN architectures fine-tuned with FFT maintain superior performance in extremely small or severely imbalanced datasets. Conversely, larger-scale FMs often under-perform in these constrained data scenarios. Nevertheless, our novel finding is that specific PEFT strategies can mitigate these challenges for FMs, offering an effective approach to adapt large pretrained models even to particularly demanding tasks such as prognosis.
The design of an effective fine-tuning strategy is a key decision point in adapting large pretrained models to prognosis tasks. Recent advances in FMs have created new opportunities for applying transfer learning in clinical prediction settings. However, the optimal fine-tuning approach remains unclear, particularly in light of growing interest in PEFT strategies. These techniques promise reduced computational and memory costs by updating only a small subset of the model’s parameters, yet their ability to retain performance under medical data constraints is still underexplored. This question is critical in the context of real clinical scenarios.
Figure 4: This figure presents one subplot for each dataset, illustrating on the $y$ axis the MCC results and on the $x$ axis the percentage of models’ parameters trained over the total. Fine tuning techniques and models have a double encoding systems, one based on colors as $\mathbf{LoRA_{r=4}}$, $\mathbf{LoRA_{r=8}}$, $\mathbf{LoRA_{r=16}}$, $\mathbf{VeRA_{r=4}}$, $\mathbf{VeRA_{r=8}}$, $\mathbf{VeRA_{r=16}}$, $\mathbf{BitFit}$, $\mathbf{IA^{3}}$, $\mathbf{LP}$, and $\mathbf{FFT}$; and one based on the model Refer to caption($CLIP\text{-}Large$), Refer to caption(MedCLIPv), Refer to caption ($BioMEdCLIP$), Refer to caption (DINOv2b), Refer to caption(DINOv2s), Refer to caption (DINOv2l), Refer to caption (MedCLIPc), and Refer to caption (PubMedCLIP).
4.2.1 Setup.
In this section, we analyze various fine-tuning techniques applied exclusively to FMs, with the combination of PEFT methods and models following Table 3. To maintain consistency, all models were trained using the same experimental settings as discussed in Section 3.5. This analysis aims to provide insights into FMs as a pretraining paradigm, rather than being tied to a specific model architecture. Figure 4 presents the results, where the x-axis represents the percentage of trained FM parameters over the total, and the y-axis indicates the average metric value obtained on the test sets.
4.2.2 Results.
We conducted a comprehensive analysis of the results presented in Figure 4, examining performance respect different characteristics of the data, including sample size and class distribution. Furthermore, we evaluated PEFT efficiency by investigating the relationship between model performance and the proportion of the model trainable parameters adapted during fine-tuning. This analysis aimed to quantify the effectiveness of PEFT in relation to their ability to achieve task-specific adaptation while modifying only a small fraction of the original model parameters.
In smaller datasets such as CAR and CC (Figure 4.c,  4.d), most fine-tuning techniques fail to deliver significant improvements, suggesting that despite their parameter efficiency, these methods struggle to adapt and generalize under data constraints. This is particularly evident in Figure 4.d, where most techniques plateau around $10\%$ MCC, with the exception of LoRAr=8 and BitFit, which perform notably better when applied to DINOv2l and DINOv2s.
For CAR, results are generally more favorable, with several experiments achieving up to $30\%$ MCC. Notably, LoRAr=4 and LoRAr=16 applied to DINOv2s reach performance levels close to $80\%$, indicating that certain configurations can still extract meaningful representations despite limited data. However, in both datasets, FFT consistently underperforms, likely due to the constraints imposed by small sample sizes. These findings suggest that the most parameter-efficient techniques are not always the most effective in low-data scenarios. Factors such as dataset complexity, class imbalance, and the pretraining knowledge of the model play a key role during fine-tuning. Some methods, like LoRA and BitFit, leverage pretrained representations better, while others, such as FFT, struggle due to excessive constraints on the dataset size. This shows that there is no direct correlation between the percentage of trainable parameters and final performance, indicating that adaptation depends on a combination of model adaptability to capture relevant features discriminant for prognosis tasks.
Another key finding observed is the impact of task imbalance on performance. This analysis is performed comparing AFC and AFCm (Figure 4.a, 4.b, as both datasets contain the same number of samples but differ in class distribution across the two tasks. In the balanced dataset AFC (Figure 4.a), all fine-tuning methods achieve results stable around $\sim 30\%$ MCC for the majority of the methods, with FFT performing best for DINOv2s and PubMedCLIP. However, in the imbalanced task AFCm (Figure 4.b), performance drops significantly despite the dataset size remaining the same. Notably, LoRA and BitFit still manage to reach $35\%$ MCC, indicating resilience to class imbalance. This highlights that dataset properties, beyond size alone, play a crucial role in fine-tuning outcomes. When fine-tuning FMs, class balance must be carefully considered, as intrinsic model biases can lead to misleadingly poor results if not properly accounted for.
A similar pattern emerges when comparing NYsmall and NYall (Figures 4.e and 4.f), which represent the same task but differ in dataset size by a factor of ten. Several fine-tuning techniques remain around $\sim 10\%$, but DINOv2s and DINOv2b achieve the best results with LoRA, BitFit, and FFT on NYall, particularly for DINOv2 and CLIP-Large. In contrast, only LoRA performs well on NYsmall. Interestingly, FFT achieves the highest performance on NYsmall, mirroring the behavior observed on AFCm. However, on NYall, its performance aligns with PEFT methods that train fewer parameters, suggesting that the larger dataset improves adaptation, reducing the need for extensive parameter updates needed in FFT. FT effectiveness is strongly influenced by the pretraining model’s inherent biases. In general, FFT achieves the best performance, particularly for PubMedCLIP and DINOv2 variants. This suggests that for efficient adaptation, a generalist model like DINOv2, despite lacking domain-specific medical pretraining, performs best across diverse tasks. In contrast, a biomedical-specific model like PubMedCLIP benefits the most from FFT, emphasizing the importance of selecting fine-tuning strategies based on both model characteristics and dataset properties.
These findings confirm that there is no one-size-fits-all fine-tuning strategy. The optimal choice depends on dataset size, class distribution, and model-specific inductive biases. While FFT remains the gold standard in large, balanced datasets, its limitations become evident in small or skewed conditions. Conversely, PEFT methods like LoRA and BitFit, though parameter-efficient, are highly sensitive to dataset properties. Given these results, in the following sections we move our attention to the evaluation of PEFT techniques under FSL conditions. These scenarios are particularly relevant in real-world clinical practice, where annotations are limited and new tasks must be adapted quickly with minimal supervision. By isolating PEFT in this extreme regime, we aim to better understand their practical utility and identify configurations that offer robust generalization despite severe training constraints.
4.3.1 Setup.
4.3.2 Results.
We analyze the results presented in Figure 5. The first key observation is that, across all datasets, performance remains substantially lower than expected. However, an interesting pattern emerges when comparing the AFC dataset in its balanced and imbalanced versions (Figures 5.a and 5.b). In AFCm, the best results were obtained with VeRA, one of the most parameter-efficient approaches, suggesting that efficiency plays a role in this specific setting. However, the overall performance across all datasets appears highly inconsistent, showing no clear evidence of a robust and repeatable trend. In AFC and NYsmall (Figure 5.e), LP achieves the best absolute performance, contrasting with the results from AFC, where LP alone performed best. This discrepancy may be linked to the parameter space, as the number of available samples is highly constrained. It is likely that the limited dataset size affects the model’s ability to effectively predict the task. Similar trends were observed in the fine-tuning settings explored in Section 4.2, where most methodologies failed to achieve satisfactory performance. Several factors could contribute to this, including the adaptability of foundation models, the efficacy of fine-tuning methods, or the complexity of the tasks themselves.
5 Conclusions
Our findings lead to several key conclusions:
  • 1.
    CNNs Remain Strong Baselines in Extreme Low-Data Regimes: CNNs, particularly when fine-tuned using FFT, consistently outperform FMs on extremely small and highly imbalanced datasets. Their lower parameter count and strong inductive biases make them less prone to overfitting and better suited for clinical scenarios with scarce labeled data.
  • 2.
    FMs with PEFT Achieve Competitive Performance on Larger Datasets: while FMs struggle in extreme low-data regimes, they demonstrate superior scalability and adaptability as dataset size increases. PEFT methods like LoRA and BitFit enable effective adaptation of FMs with minimal parameter updates, achieving competitive results with significantly lower computational overhead compared to FFT.
  • 3.
    PEFT Methods Are Highly Sensitive to Class Imbalance: across all experiments, the performance of PEFT techniques deteriorates under severe class imbalance, particularly when minority classes represent critical clinical outcomes such as ICU admission or mortality. While FFT remains more robust under these conditions, PEFT methods can still provide substantial gains when dataset balance is moderately improved.
  • 4.
    Few-Shot Learning Remains Challenging, but Selective PEFT Strategies Offer Promise: in FSL scenarios, PEFT methods applied to smaller FM variants, such as DINOv2s, demonstrate an ability to extract meaningful representations even with extremely limited training data. However, the success of these approaches strongly depends on careful model selection and hyperparameter tuning, highlighting the need for further research on PEFT optimization in clinical few-shot settings.
  • 5.
    No Universally Optimal Fine-Tuning Strategy: our benchmark highlights that the effectiveness of a fine-tuning approach depends on the interplay between model architecture, dataset characteristics, and computational constraints. CNNs remain a reliable option for low-resource settings, while FMs paired with appropriate PEFT methods offer scalable and efficient solutions as data availability increases.
Our analysis provides actionable guidelines for selecting fine-tuning strategies in real-world medical prognosis applications. It also underscores the need for continued exploration of PEFT techniques, especially in the context of Few-Shot Learning, to fully unlock the potential of FMs in data-constrained and high-impact clinical environments.
Limitations.
This study focuses exclusively on COVID-19 chest X-ray prognosis tasks, which, while clinically relevant, may not generalize to other imaging modalities or diseases. Additionally, we rely on publicly available datasets that differ in annotation quality and may not fully capture real-world clinical heterogeneity. Although we evaluated a broad range of models and fine-tuning strategies, the search space remains constrained in terms of PEFT configurations and hyperparameters. Nevertheless, the breadth and standardization of our benchmark enable a consistent and informative assessment of fine-tuning strategies across heterogeneous prognostic tasks, while leaving the exploration of medical prognosis across diverse imaging modalities to future work.
5.0.1 Acknoledgements
Appendix A Overall Benchmark Results
@@
Table 4: This table presents the $\mathbf{MCC}$ scores for all the experiments divided in two main sections by a double line, to discriminate the results obtained from CNNs (top) and FMs (bottom). Each row displays a model’s performance across multiple tasks under different fine-tuning (FT) approaches, with the best fine-tuning approach per task highlighted in bold. Cells highlighted in green and orange indicate the top and second-best results, respectively, for each dataset task across all model and fine-tuning combinations.
$\mathbf{CC}$1.4 $\pm$ 14.8 1.1$\pm$ 9.2 2.1 $\pm$ 7.2 - - - 6.5 $\pm$ 0.8 - 3.0 $\pm$ 12.3 46.8 $\pm$ 25.2
CC 2.2 $\pm$ 19.3 -2.7 $\pm$ 11.5 8.9 $\pm$ 16.0 - - - 6.5 $\pm$ 0.8 - 3.4 $\pm$ 20.9 3.7 $\pm$ 3.8
CC -3.4 $\pm$ 10.9 -0.1 $\pm$ 8.7 2.3 $\pm$ 15.1 - - - 8.0 $\pm$ 2.9 - -1.5 $\pm$ 9.4 32.5 $\pm$ 44.0
CC 1.6 $\pm$ 10.9 1.5 $\pm$ 12.0 3.1 $\pm$ 12.4 - - - 4.0 $\pm$ 16.8 - 7.4 $\pm$ 2.4 8.5 $\pm$ 4.3
CC -4.7 $\pm$ 21.4 -4.7 $\pm$ 21.4 -4.7 $\pm$ 21.4 -3.3 $\pm$ 18.1 -3.3 $\pm$ 18.1 -3.3 $\pm$ 18.1 3.2 $\pm$ 12.1 -0.2 $\pm$ 15.2 -0.8 $\pm$ 9.3 6.8 $\pm$ 1.3
CC -6.4 $\pm$ 15.1 -4.6 $\pm$ 10.7 -7.5 $\pm$ 10.4 -7.1 $\pm$ 6.2 -1.0 $\pm$ 13.7 -1.1 $\pm$ 13.0 3.5 $\pm$ 6.4 3.3 $\pm$ 12.6 2.6 $\pm$ 19.8 6.5 $\pm$ 0.8
CC -2.6 $\pm$ 8.6 37.7 $\pm$ 27.6 -3.5 $\pm$ 11.3 1.1 $\pm$ 11.1 8.1 $\pm$ 5.6 7.2 $\pm$ 6.8 39.2 $\pm$ 23.8 0.9 $\pm$ 11.8 1.8 $\pm$ 5.6 2.2 $\pm$ 9.3
CC 8.3 $\pm$ 5.7 6.5 $\pm$ 0.8 10.8 $\pm$ 16.0 10.1 $\pm$ 5.5 11.0 $\pm$ 5.9 11.0 $\pm$ 5.9 6.8 $\pm$ 6.7 4.0 $\pm$ 10.8 15.5 $\pm$ 12.9 4.3 $\pm$ 5.5
CC 7.7 $\pm$ 2.3 6.1 $\pm$ 0.9 1.4 $\pm$ 7.2 - - - 5.1 $\pm$ 10.3 -6.7 $\pm$ 8.3 1.5 $\pm$ 7.0 -3.6 $\pm$ 9.2
CC -0.3 $\pm$ 7.8 1.2 $\pm$ 6.2 1.2 $\pm$ 6.2 1.2 $\pm$ 6.2 1.2 $\pm$ 6.2 1.2 $\pm$ 6.2 6.5 $\pm$ 0.8 -7.1 $\pm$ 10.3 7.6 $\pm$ 11.6 17.8 $\pm$ 16.4
$\mathbf{NY_{small}}$0.6 $\pm$ 10.7 1.4 $\pm$ 10.5 1.4 $\pm$ 10.5 1.6 $\pm$ 10.6 1.4 $\pm$ 10.5 1.6 $\pm$ 10.6 13.2 $\pm$ 0.0 2.0 $\pm$ 5.8 21.1 $\pm$ 9.0 65.0 $\pm$ 25.2

### v2

@@ 4 Results and Discussion
Figure 2: Distribution of performance scores across fine-tuning methods and datasets. Each box plot summarizes the mean performance of all models fine-tuned with a given technique on a specific dataset. Boxes indicate the inter-quartile range for all the mean result obtained from all the models fine-tuned with a specific technique on a specific dataset (divided by color), while the central line represents the median and whiskers extend to the minimum and maximum values.
This section presents and discusses the main experimental results obtained following the experimental methodology described in Section 3. All reported results are evaluated using the MCC and PR-AUC scores, which serve as the primary performance metrics. A complete summary is presented in Table 4 and Table 5, which report all experimental outcomes and highlight the top- and second-performing model–technique pairs for each dataset, determined according to the highest mean value of the MCC and PR-AUC metrics, respectively. From the MCC score in Table 4, we further derived Figure 2, displaying the inter-quartile range of the mean results obtained from models fine-tuned with each specific technique on each dataset.
In addition, to formally support comparative claims, all methods were further assessed through a non-parametric statistical analysis. Specifically, we employed the one-tailed Wilcoxon signed-rank test on the rank distributions derived from the mean MCC values of the models across all datasets. This approach yields a pairwise significance matrix (shown in Table 6 and Table 7), which reports, for each method pair, the number of datasets in which the method on a given row outperforms the method on the corresponding column and the eventual statistical significance in terms of $p$-values, annotated using the conventional significance levels “*” for $p<0.05$. This non-parametric framework provides a reliable indication of whether observed differences are statistically meaningful rather than arising from random variation.
The following results are organized into two main parts. The Setup paragraph outlines the specific experimental configuration, including the models and PEFT methods analyzed, while the Results paragraph provides an in-depth analysis of the results obtained through the combination of models, datasets, and fine-tuning strategies described in Section 3.
[…]
$C$ 2.2 $\pm$ 19.3 -2.7 $\pm$ 11.5 8.9 $\pm$ 16.0 - - - 6.5 $\pm$ 0.8 - 3.4 $\pm$ 20.9 3.7 $\pm$ 3.8
$C$ -3.4 $\pm$ 10.9 -0.1 $\pm$ 8.7 2.3 $\pm$ 15.1 - - - 8.0 $\pm$ 2.9 - -1.5 $\pm$ 9.4 32.5 $\pm$ 44.0
$C$ 1.6 $\pm$ 10.9 1.5 $\pm$ 12.0 3.1 $\pm$ 12.4 - - - 4.0 $\pm$ 16.8 - 7.4 $\pm$ 2.4 8.5 $\pm$ 4.3
$C$ -4.7 $\pm$ 21.4 -4.7 $\pm$ 21.4 -4.7 $\pm$ 21.4 -3.3 $\pm$ 18.1 -3.3 $\pm$ 18.1 -3.3 $\pm$ 18.1 3.2 $\pm$ 12.1 -0.2 $\pm$ 15.2 -0.8 $\pm$ 9.3 6.8 $\pm$ 1.3
$C$ -6.4 $\pm$ 15.1 -4.6 $\pm$ 10.7 -7.5 $\pm$ 10.4 -7.1 $\pm$ 6.2 -1.0 $\pm$ 13.7 -1.1 $\pm$ 13.0 3.5 $\pm$ 6.4 3.3 $\pm$ 12.6 2.6 $\pm$ 19.8 6.5 $\pm$ 0.8
$C$ -2.6 $\pm$ 8.6 37.7 $\pm$ 27.6 -3.5 $\pm$ 11.3 1.1 $\pm$ 11.1 8.1 $\pm$ 5.6 7.2 $\pm$ 6.8 39.2 $\pm$ 23.8 0.9 $\pm$ 11.8 1.8 $\pm$ 5.6 2.2 $\pm$ 9.3
$C$ 8.3 $\pm$ 5.7 6.5 $\pm$ 0.8 10.8 $\pm$ 16.0 10.1 $\pm$ 5.5 11.0 $\pm$ 5.9 11.0 $\pm$ 5.9 6.8 $\pm$ 6.7 4.0 $\pm$ 10.8 15.5 $\pm$ 12.9 4.3 $\pm$ 5.5
$C$ 7.7 $\pm$ 2.3 6.1 $\pm$ 0.9 1.4 $\pm$ 7.2 - - - 5.1 $\pm$ 10.3 -6.7 $\pm$ 8.3 1.5 $\pm$ 7.0 -3.6 $\pm$ 9.2
$C$ -0.3 $\pm$ 7.8 1.2 $\pm$ 6.2 1.2 $\pm$ 6.2 1.2 $\pm$ 6.2 1.2 $\pm$ 6.2 1.2 $\pm$ 6.2 6.5 $\pm$ 0.8 -7.1 $\pm$ 10.3 7.6 $\pm$ 11.6 17.8 $\pm$ 16.4
$\mathbf{NY_{small}}$ -0.6 $\pm$ 10.7 1.4 $\pm$ 10.5 1.4 $\pm$ 10.5 1.6 $\pm$ 10.6 1.4 $\pm$ 10.5 1.6 $\pm$ 10.6 13.2 $\pm$ 0.0 2.0 $\pm$ 5.8 21.1 $\pm$ 9.0 65.0 $\pm$ 25.2
Table 4: This table presents the $\mathbf{MCC}$ scores for all the experiments divided in two main sections by a double line, to discriminate the results obtained from CNNs (top) and FMs (bottom). Each row displays a model’s performance across multiple tasks under different fine-tuning approaches, with the best fine-tuning approach per task highlighted in bold. Cells highlighted in green and orange indicate the top and second-best results, respectively, for each dataset task across all model and fine-tuning combinations.